Find three consecutive integers who's sum is 171
Answer by sudhanshu_kmr(1152)

LET NUMBERS ARE X ,X+1, X+2


X + X+1 + X+2 =171

=> 3X +3 = 171

=> X = (171-3)/ 3 = 56

NUMBERS ARE 56, 57, 58